What mass of water was used if the temperature went from 55oc to 100oc and the energy used was 65,000 j?
1
Expert's answer
2020-05-24T03:09:57-0400

Solution:

The amount of thermal energy stored or released as the temperature of a system changes can be calculated using the equation:

ΔEtQ = m × c × ΔT

where:

  • ΔEtQ - change in thermal energy (J)
  • m - mass (g)
  • c - specific heat capacity (J/g°C)
  • ΔT - temperature change (°C)


ΔEtQ = 65000 J

m = unknown

c(H2O) = 4.184 J/g°C

ΔT = Tfinal - Tinitial = (100 - 55)°C = 45°C


m = ΔEtQ / (c × ΔT)

m = (65000 J) / (4.184 J/g°C × 45°C) = 345.23 g

m(H2O) = 345.23 g


Answer: 345.23 g of water (H2O) was used.